Use the fundamental theorem of homomorphism for groups to prove the following theorem, which is called the zassenhaus


1
Expert's answer
2021-05-14T05:49:36-0400

Zassenhaus's Lemma:


Statement:

Let G be a group; let "H,H' ,K ,K'"  be subgroups of G such that H' is a normal subgroup of H and K' is a normal subgroup of K. Then "H'\cdot (H\cap K')"  is a normal subgroup of "H'\cdot(H\cap K)" ; likewise, "K'\cdot (K\cap H')"  is a normal subgroup of "K'\cdot(H\cap K)"  ; furthermore, the quotient groups


"(H'\cdot (H\cap K))/(H'\cdot (H\cap K'))"

and


"(K'\cdot(H\cap K))/(K'\cdot (K\cap H'))"

are isomorphic.


Proof:

We first note that "H\cap K"  is a subgroup of H. Let "\eta"  be the canonical homomorphism from H to H/H'. Then "(\eta^{-1}\circ\eta)(H\cap K)=H'\cdot (H\cap K)" , so this indeed a group. Also, note that "H\cap K'"  is a normal subgroup of "H\cap K" . Hence


"(\eta^{-1}\circ\eta)(H\cap K')=H'\cdot (H\cap K')"

is a normal subgroup of


"(\eta^{-1}\circ\eta)(H\cap K)=H'\cdot (H\cap K)"

Now,

Let "\lambda" be the canonical homomorphism from "H'\cdot(H\cap K)"  to "(H'\cdot(H\cap K))/(H'\cdot (H\cap K'))" . Now, note that


"(H\cap K )\cap(H'\cdot (H\cap K'))=(H'\cap K)\cdot (H\cap K')"



Thus by the group homomorphism theorems, groups "(H'\cdot (H\cap K))/(H'\cdot (H\cap K'))"  and "(H\cap K)/((H'\cap K)\cdot(H\cap K'))" are isomorphic. The lemma then follows from symmetry between H and K.